OpenStudy (anonymous):

can someone please explain how you would find the exact value of sqrt(2)/sqrt(8)

OpenStudy (anonymous):

\[\sqrt{2}/2\sqrt{2}\]

OpenStudy (anonymous):

= 1/2 = 0.5

OpenStudy (anonymous):

8 = 2*2*2

OpenStudy (ash2326):

@lala1234 do you understand this?